OpenStudy (anonymous):

1/4 + 1/5= Type a whole number or a simplified fraction.

OpenStudy (anonymous):

put the fractions over common denominator...and then add common denominator is 20 to me

OpenStudy (anonymous):

so 1/20 + 1/20

OpenStudy (anonymous):

to put over common denomitor - what you do to one part of fraction must be done to the other part of the fraction 1/4 to get this over 20 both top and bottom must be multiplied by 5 try again...

OpenStudy (anonymous):

5/20 right+ 1/5

OpenStudy (anonymous):

yes do the same for the 1/5 fraction...what must you multiply top and bottom by to get 20 on the bottom

OpenStudy (anonymous):

5/25

OpenStudy (anonymous):

you trying to get 20 in the bottom not 25 almost

OpenStudy (anonymous):

so how would i get it if i multiplied by 5

OpenStudy (anonymous):

no, 5 goes to 20 by multiply 4

OpenStudy (anonymous):

4/20 right?

OpenStudy (anonymous):

yes now add both of your new fractions together

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }+\frac{ 1 }{ 5 }=\frac{ 5 }{ 20 }+\frac{ 4 }{ 20 }=\frac{ 9 }{ 20 }\]

OpenStudy (anonymous):

9/20

OpenStudy (anonymous):

perfect thanks for working it out yourself in order to understand the problem Good job

OpenStudy (anonymous):

Thank you for helping me